OpenStudy (anonymous):

4(x-b)=x SHOW WORK PLEASE

OpenStudy (anonymous):

first you should distribute that 4 on the left side

OpenStudy (anonymous):

If it is solve for x 4(x-b)=x 4x-4b=x -4b=-3x